OpenStudy (anonymous):

Determine the solution set of (x - 3)^2 = 49.

OpenStudy (anonymous):

{6, -8} {4, 10} {10, -4}

Parth (parthkohli):

Find the square root of both sides first.

OpenStudy (anonymous):

(x-3)^2=\[x-3=\pm \sqrt{49}\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow x - 3 = 7 }\) or \( \color{Black}{x - 3 = -7 }\) This is because both +7 and -7 are square roots of 49.

OpenStudy (anonymous):

According to the square root theorem

OpenStudy (anonymous):

So it's the latter

OpenStudy (anonymous):

ok

Parth (parthkohli):

@purplec16 no need to answer more :)

OpenStudy (anonymous):

oh, lol...

Parth (parthkohli):

Tyler, solve both equations and see both the answers. Tally.

OpenStudy (anonymous):

oky doky

OpenStudy (anonymous):

Can you help me with my question parth?

OpenStudy (anonymous):

Please :)

OpenStudy (anonymous):

what equation do i solve

Parth (parthkohli):

I gave you the equations. @purplec16 SURE :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

10,-4 thanks parth